OpenStudy (anonymous):

Write a line parallel to y = 2x+4 and passing through the point (3,8)

OpenStudy (anonymous):

parallel lines have the same slope, so we're looking at a slope of 2 here...so we have y=2x+b.....so no, we're going to fill in our x and y value from the point. we get.... 8 = 2(3) + b 8=6 + b 2 = b. So, the final equation would be y = 2x + 2